# Model-free reinforcement learning with noisy actions for automated experimental control in optics

### v2

@@ 5 Experimental results in the optics lab
The agent was trained in our lab on components detailed in Section 3.1. Our training speed is limited by the time the actuators take to move, leading to each environment step taking about a second. We let each training run until its return starts to converge. For $P_{\text{goal}}\leq 0.87$, this took around 20 hours or $4\cdot 10^{4}$ steps. That we can train successfully in only $4\cdot 10^{4}$ steps is a result of the environment shaping discussed in Appendix C. If we set a very high goal power ($P_{\text{goal}}=0.9$), the training takes much longer ($2\cdot 10^{5}$ time steps, which sums up to nearly 4 days of training).
[…]
In the different training runs, we changed algorithms and goal powers. For goal powers between $P_{\text{goal}}=0.85$ and $P_{\text{goal}}=0.87$, training runs start to converge at around $4\cdot 10^{4}$ steps. For higher goal powers, e.g., $P_{\text{goal}}=0.9$, this is not the case anymore. This is shown in Figure 2 (a). We tested both SAC and TQC by performing two training runs per algorithm on the experiment with a goal power of $P_{\text{goal}}=0.85$. We can see that in the beginning, the return rises quicker for SAC but is slightly outperformed by TQC later on. This is why we chose TQC for all other experiments presented here.
@@
When we choose $P_{\text{goal}}=0.9$, we need significantly more training steps ($\geq 2\cdot 10^{5}$). Because of that, as discussed before, we also tested pre-training on lower goals on the experiment, i.e., we started the training with a lower goal power $P_{\text{goal}}=0.85$ and increased it in small increments to $P_{\text{goal}}=0.9$ over the course of the first $10^{5}$ training steps and left it at that for the next $10^{5}$ steps. These training runs are shown in Figure 2 (b). The normalized return of the agent first pre-trained on lower powers always drops after changing to a higher goal power, as it first needs to learn to handle the conditions of the changed environment. We can also see that the normalized return reaches lower values the higher the goal gets. This is expected as the task gets harder each time. The normalized return for the agent pre-trained on lower goals reaches higher values than the one starting from scratch. Additionally, we can use intermediate agents to align the experiment to lower powers. We conclude that curriculum learning was helpful for such a high goal in our experiment (see also [76]). Although we also found pre-training on the virtual testbed with an added noise model helpful (see Appendix D), we focus on in-situ training, as we want to find strategies that can work on experiments where a noise model is hard or impossible to obtain.
[…]
@@
Table 3: Reset by hand vs. automatic reset. We tested the agent pre-trained on lower goals with $P_{\text{goal}}=0.9$ as described in Figure 2 (automated reset) or resetting by tilting the hand-steering motors $29$ times. We compare the mean of the power at the start $\bar{P}_{0}$ and end of the episode $\bar{P}_{T}$, the empirical probability of reaching the goal $p[\text{goal}]$ or failing $p[\text{fail}]$, and mean of the number of environment steps needed to reach the goal $\bar{\tau}_{\text{goal}}$.
$\bar{P}_{0}$ $\bar{P}_{T}$ $p[\text{goal}]$ $p[\text{fail}]$ $\bar{\tau}_{\text{goal}}$
Reset by hand $0.384$ $0.91411$ $0.86$ $0.03$ $8.333$
Automatic reset $0.465$ $0.909936$ $0.9$ $0.03$ $8.22$
@@ 6 Summary and Outlook
We have shown that our model-free RL agent successfully learns to couple laser light into an optical fiber, reaching the same efficiencies as a human expert while generally being faster. We find that sample-efficient algorithms that use a replay buffer, such as SAC and TQC, are a must to overcome the challenge of otherwise not manageable training time. Partial observability can be dealt with by carefully tuning the observations. Furthermore, our study suggests that curriculum learning can help to achieve more difficult goals.
As we train directly on the experiment, the agent learns to handle the specific noise present, and we can avoid creating an accurate simulation of the task. This makes our method suitable for setups where it is impossible to model the noise accurately.
@@
Further exploration could include investigations on how the agents perform for other transverse optical modes of light, including multiple local maxima, and investigating the performance of model-based or hybrid algorithms. We start our RL training procedure under conditions where there is low power. This raises the question how, or with how many additional sensors, we could generalize this to the case of starting with no power. Furthermore, it might be interesting to explore replacing the use of a history as our observation by PID-inspired RL [77] and to investigate whether pretraining on expert demonstrations could speed up the learning process [78, 79, 80].
Our experiments are a first step towards the extensive use of RL in our quantum optics laboratory. Optical experiments typically require various control loops to stabilize the experimental degrees of freedom against perturbations. These locks significantly increase the complexity of the experiments. For example, maintaining the length of optical cavities to achieve resonant light field enhancement requires complex components such as phase modulators [81], homodyne detectors [82, 83], or split detectors [84, 85]. RL offers the potential for streamlined control loops that rely solely on the measurement of power in the reflection and transmission of the resonator. This could enable novel control strategies, such as phase control of squeezed vacuum states. These states are characterized by unique quantum noise properties but are otherwise dark. Consequently, phase control typically requires an auxiliary laser field [86] or introduces unwanted phase noise [87]. RL has the potential to provide a noise-free solution without the need for additional laser fields, which is particularly relevant for large-scale on-chip squeezing experiments in the field of quantum information [88].
@@
Acknowledgments and Disclosure of Funding
We thank Amir Abolfazli, Timko Dubielzig, Ludwig Krinner, and Kilian Stahl for helpful discussions, Pratik Chakraborty and Manuel Schimanski for help in the lab, and Nived Johny, Jonas Junker, Arindam Saha and Bernd Schulte for both of them. LR thanks Mika Nina Seibicke and Alina Jankowski for lab sitting and emotional support. This work was supported by the Quantum Valley Lower Saxony (QVLS), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through SFB 1227 (DQ-mat), Germany’s Excellence Strategy EXC-2123 QuantumFrontiers 390837967, and Excellence Strategy EXC-2122 PhoenixD 390833453, the BMBF projects ATIQ and QuBRA, the BMWK project ProvideQ, and the Australian Research Council Centre of Excellence for Quantum Computation and Communication Technology (project number CE170100012).
@@ Appendix A Additional details about the experimental setup
[…]
For training on the experiment, we used an NVIDIA GeForce RTX 4070 GPU. In addition to the usual packages [89, 90, 91, 92], we used PyLabLib, Thorlabs Kinesis, PyVisa, Keysight Connection Expert and safe-exit [93, 94, 95, 96, 97] for communication with the experiment.
@@ Appendix B Explanation of the imprecision in the mirror steering motors and its characterization
Backlash is a phenomenon that is present when a load is not directly connected to a motor, such as in geared mechanical systems [98]. Dependent on the exact geometry of the system, i.e., mechanical tolerances, amount of gears, etc., it may resemble hysteresis between the expected and actual position or a dead-zone, where moving the actuator has no effect on the actual position whenever the rotational direction is reversed. It is thus hard to model and predict a priori. Control has to be implemented based on the specific system and its use. These control schemes include hysteresis models, dead-zone models, and PI control. However, additional sensors are needed to get accurate feedback if multiple actuators are used. Step loss results from the difference in static and dynamic torque of a motor. The motor steps result in a linear actuation, which changes the tilt of the mirror mount. Different tilt angles lead to different static loading of the motor and gearbox. This may lead to the initial step(s) being lost, as the motor can not deliver the starting torque, resulting in a partial step. Without feedback from, e.g., an encoder, this leads to a difference between the expected and actual position. Lastly, the non-orthogonal degrees of freedom are a result of the kinematic mirror mounts used and their mounting. Usually, this error is small for well-designed kinematic mirror mounts.
@@ Appendix C Environment and agent tuning on virtual testbed
$(\displaystyle P(x_{m1},y_{m1},x_{m2},y_{m2})=0.92\mathrm{exp}\Bigg{(}-\frac{1}{2}\Bigg{(}$$\left(\frac{x_{m1}-5470785}{11994}\right)^{2}+\left(\frac{y_{m1}-5573194}{19145}\right)^{2}$
@@ C.1 Reward Hyperparameters
×$\times\{1,5,10\}$$\times\{1,5\}^{2}$$\times\{1,5\}^{2}$
@@ C.2 Episode Length
We want to find a good trade-off between reaching the goal quickly and reaching it reliably. Using the same parameters as for reward shaping, we tested different episode lengths, in particular, $T=5,10,...,50$. The results are shown in Figure 7.
@@ C.5 Observation
We tested history lengths of $n=1,...,6$ and the maximum sensible length $n=T=30$ with $P_{\text{goal}}=0.85$. The results can be found in Figure 10 (a).
Figure 10: Comparison of different observations and maximum actions. (a)-(c) show the normalized return in dependence of the training step using the parameters in Tables 1 and 2, TQC, and $P_{\text{goal}}=0.85$ for different history lengths $n$ in (a), leaving out different parts of the observation in (b), and with or without including the absolute positions in the observation in (c). The models with and without the absolute positions were then tested $100$ times each in an environment in which $k=0,...,4$ means of the underlying Gaussian $\mu_{i}$ were shifted by ±$\pm\sigma_{i}$, i.e., $\mu_{i}^{\prime}=\mu_{i}\pm\sigma_{i}$. (d) shows the probability of reaching the goal against the number of shifts $k$. The plots show the mean with $2\sigma$ error bands/bars created by multiple runs and, in (a)-(c), smoothing.
@@
Additionally, we tested how agents perform in an environment that includes the absolute position of the actuators in the observation compared to one that does not. The normalized return against the training step can be found for both configurations, using the parameters in Tables 1 and 2, TQC, and $P_{\text{goal}}=0.85$, in Figure 10 (c). As expected, the agent learns faster and reaches a higher normalized return if those absolute positions are included. However, the main application for our agent is to recouple the light into the fiber after other parts of the experiment have been misaligned. That means, that the optimal positions, i.e., the means $\mu_{i}$ of the underlying Gaussian, change, and the agent still has to be able to reach the goal. Hence, we tested the agent $100$ times each in environments in which $k=0,...,4$ means of the underlying Gaussian $\mu_{i}$ were shifted by ±$\pm\sigma_{i}$, i.e., $\mu_{i}^{\prime}=\mu_{i}\pm\sigma_{i}$. Figure 10 (d) shows the probability of reaching the goal against $k$. If no shifts occur, the agent with the absolute position as part of the observation performs slightly better than the one without. However, this quickly changes as more shifts are applied. The agent with absolute position performs much worse if any shifts appear. On the other hand, the agent not observing the absolute position performs well independent of shifts.

### v3

@@ 5 Experimental results in the optics lab
The agent was trained in our lab on components detailed in Section 3.1. Our training speed is limited by the time the actuators take to move, leading to each environment step taking about a second. We train each agent until the return starts to converge as depicted in Figure 2 (a)-(b). For $P_{\text{goal}}\leq 0.87$, this took around 20 hours or $4\cdot 10^{4}$ steps. That we can train successfully in only $4\cdot 10^{4}$ steps is a result of the environment shaping discussed in Appendix C. If we set a very high goal power ($P_{\text{goal}}=0.9$), the training takes longer: for CrossQ and SAC about $1.2\cdot 10^{5}$ time steps, which corresponds to 2 days, and for TQC $2\cdot 10^{5}$ time steps, which sums up to nearly 4 days of training.
Using TQC, SAC, and CrossQ, we performed two training runs per algorithm on the experiment with a goal power of $P_{\text{goal}}=0.85$. In Figure 2 (a), we can see that in the beginning, the return rises quicker for SAC and CrossQ but is slightly outperformed by TQC later on. However, these differences are small.
@@
When we choose $P_{\text{goal}}=0.9$, as shown in Figure 2 (b), we need significantly more training steps. Because of that, as discussed before, we also investigated pre-training using TQC. Although we found pre-training on the virtual testbed with an added noise model helpful (see Appendix D), we focus on in-situ training, as we want to find strategies that can work on experiments where a noise model is hard or impossible to obtain. Thus, we implemented pre-training on lower goals directly on the experiment using TQC (light blue), i.e., we started the training with the lower goal power $P_{\text{goal}}=0.85$, increased it in small increments to $P_{\text{goal}}=0.9$ over the course of the first $10^{5}$ training steps and left it at that for the next $10^{5}$ steps. The normalized return of the agent first pre-trained on lower powers always drops after changing to a higher goal power, as it first needs to learn to handle the conditions of the changed environment. We can also see that the normalized return reaches lower values the higher the goal gets. This is expected as the task gets harder each time. For TQC, the normalized return for the agent pre-trained on lower goals reaches higher values than the one starting from scratch (green). Additionally, we can use intermediate agents to align the experiment to lower powers.
The return for SAC (dark pink) reaches return values similar to those of TQC without pre-training. Both are surpassed by CrossQ (brown) and the pre-trained TQC agent, which are again similar. The training time is shorter for SAC and CrossQ than for the respective TQC agents. We conclude that in our experiment curriculum learning was helpful for such a high goal when using TQC (see also Stooke et al. (2021)) while this is not needed when employing CrossQ.
Figure 2: Experimental results. (a) and (b) show the normalized return plotted against training steps for different algorithms and goal powers. In (a), we use TQC (blue), SAC (pink), and CrossQ (orange) with $P_{\text{goal}}=0.85$ and TQC with $P_{\text{goal}}=0.87$ (yellow). Of the agents trained on $P_{\text{goal}}=0.9$, shown in (b), one TQC agent, shown in light blue, was first pre-trained on lower, over time increasing, goal powers (0.85, 0.875, 0.89, 0.90.85,\,0.875,\,0.89,\,0.9). The others, i.e., TQC (green), SAC (dark pink), and CrossQ (brown), were trained directly with $P_{\text{goal}}=0.9$. (c)-(e) show our results when testing the agents marked with a star in (a) and (b). For testing, we reset and let the agents fiber couple, which is repeated $100$ times for the RL agent and $25$ times for the human expert. If the RL agents do not reach their goal within 30 steps, we reset the environment (still measuring the time), and the agent continues from there. This is repeated until the agent reaches its goal. (c) shows the power plotted against the time for the first $30\text{\,}\mathrm{s}$. The error band is clipped to the power range $[0,1]$. (d) shows the time each agent trained to $P_{\text{goal}}=0.9$ and a human expert took to reach that goal. (e) compares the number of environment steps it took for each of the RL agents tested to reach their goal after their last reset. (a)-(c) show the (smoothed) mean with $2\sigma$ error bands created by smoothing and/or multiple runs. The error bands in (c) additionally include a power measurement error of $2\%$.
To understand the help for our everyday lab work, we tested different agents in fiber coupling (marked with a star in Figure 2 (a) and (b)). The test results are shown in Figure 2 (c)-(e). Each of the RL agents was tested a hundred times. We measure the power over time. If the agent does not reach the goal during an episode, we reset the environment, and the agent tries from there. One episode was at most $40\text{\,}\mathrm{s}$, and the longest attempt took around $350\text{\,}\mathrm{s}$. Figure 2 (c) shows how the averaged coupled power develops in the first $30\text{\,}\mathrm{s}$. Generally, the power rises quicker in the first seconds for agents that try to reach only $P_{\text{goal}}=0.85$. Interestingly, even though CrossQ in total is faster in achieving the fiber coupling goal, for the first few seconds, the pre-trained TQC agent has the steepest power rise of all the agents trained with $P_{\text{goal}}=0.9$.
Figure 2 (e) shows the number of steps it took for the agents to reach their goal after their last reset, i.e., in the successful episode. In most cases, this corresponds to the first episode, and the agents need no additional resets. More precisely, the empirical probability of successful coupling in the first episode ranges from $95\%$ (SAC) to $98\%$ (CrossQ) for $P_{\text{goal}}=0.85$ and $83\%$ (TQC without pre-training) to $90\%$ (TQC pre-trained) for $P_{\text{goal}}=0.9$. Of the agents trained with $P_{\text{goal}}=0.85$, the CrossQ and TQC agents behave very similarly, taking only five steps at most in $75\%$ of the cases. The corresponding SAC agent is only half as fast and performs similarly to the TQC agent trained with $P_{\text{goal}}=0.87$. For $P_{\text{goal}}=0.9$, the CrossQ agent clearly outperforms the other agents, taking in $75\%$ of the cases at most seven steps. Of the remaining $P_{\text{goal}}=0.9$ agents, the pre-trained TQC agent performs best.
In Figure 2 (d), we show the total time it took each agent to reach a goal of $P_{\text{goal}}=0.9$. The results are similar to the results from (e) in that CrossQ is fastest, followed by TQC with pre-training, TQC without pre-training, and lastly SAC. For comparison, we also tested a human expert $25$ times on how long they would take to couple the fiber to $P_{\text{goal}}=0.9$ using the hand steering mirrors. This, however, is not a fair comparison. The RL agents can change all four degrees of freedom at once. The experimenter, on the other hand, has access to more information, e.g., the continuous power measurement while moving an actuator, and does not have to deal with the imprecision in the actuators, which means they can easily go back to an observed maximum. Despite this, we can see that the RL agents are generally faster, but all besides the CrossQ agent take longer in a few cases, where the agent needs several episodes to get to the goal. The CrossQ agent reaches the goal faster than the human expert in $100\%$ of the cases. In conclusion, we show that by using RL, we can consistently couple light into an optical fiber to high efficiencies, despite the noisy actions.
@@
Table 3: Reset by hand vs. automatic reset. We tested the agent ($P_{\text{goal}}=0.9$, pre-trained on lower goals as described in Figure 2 (b)) using two types of resets: using the reset procedure described in Section 4 (automatic reset using the motors) or resetting by tilting the hand-steering mirrors, modeling arbitrary experimental perturbation, $29$ times. We compare the mean of the power at the start $\bar{P}_{0}$ and end of the episode $\bar{P}_{T}$, the empirical probability $p[\text{goal}]$ of reaching the goal in the first episode, and mean of the number of environment steps needed to reach the goal $\bar{\tau}_{\text{goal}}$.
$\bar{P}_{0}$ $\bar{P}_{T}$ $p[\text{goal}]$ $\bar{\tau}_{\text{goal}}$
Reset by hand $0.384$ $0.914$ $0.86$ $8.33$
Automatic reset $0.465$ $0.910$ $0.9$ $8.22$
@@ 6 Summary and Outlook
We have shown that our model-free RL agents successfully learn to couple laser light into an optical fiber, reaching the same efficiencies as a human expert while generally being faster. We find that sample-efficient algorithms that use a replay buffer, such as CrossQ, SAC, and TQC, are a must to overcome the challenge of otherwise not manageable training time. The use of CrossQ has a clear advantage. It halves the training time with respect to TQC. Additionally, the trained agent reaches the goal faster than the agents trained with the other two algorithms. Furthermore, our study suggests that curriculum learning can help to achieve more difficult goals.
As we train directly on the experiment, the agent learns to handle the specific noise present, and we can avoid creating an accurate simulation of the task. This makes our method suitable for setups where it is impossible to model the noise accurately. Partial observability can be dealt with by carefully tuning the observations.
@@
Further exploration could include investigations into how the agents perform for other transverse optical modes of light, including those with multiple local maxima. Secondly, we start our RL training procedure under low-power conditions. This raises the question of how, or with how many additional sensors, we could generalize this to the case of starting with no power. Furthermore, to reduce training time, investigating the performance of model-based or hybrid algorithms and other methods to increase sample efficiency is highly relevant. In addition, it might be interesting to explore replacing the use of a history as our observation by PID-inspired RL (Char & Schneider, 2023) and to investigate whether pre-training on expert demonstrations (Vecerik et al., 2018; Mart´ınez et al., 2017; Ram´ırez et al., 2022) as well as a modified procedure of pre-training on the virtual testbed (see Appendix D.3) could speed up the learning process. However, in this study we focus on showcasing direct training on the real experiment.
Our experiments are a first step towards the extensive use of RL in our quantum optics laboratory. Optical experiments typically require various control loops to stabilize the experimental degrees of freedom against perturbations. These locks significantly increase the complexity of the experiments. For example, maintaining the length of optical cavities to achieve resonant light field enhancement requires complex components such as phase modulators (Drever et al., 1983), homodyne detectors (Hansch & Couillaud, 1980; Heurs et al., 2009), or split detectors (Shaddock et al., 1999; Chabbra et al., 2021). RL offers the potential for streamlined control loops that rely solely on the measurement of power in the reflection and transmission of the resonator. This could enable novel control strategies, such as phase control of squeezed vacuum states. These states are characterized by unique quantum noise properties but are otherwise dark. Consequently, phase control typically requires an auxiliary laser field (Vahlbruch et al., 2006) or introduces unwanted phase noise (McKenzie et al., 2005). RL has the potential to provide a noise-free solution without the need for additional laser fields, which is particularly relevant for large-scale on-chip squeezing experiments in the field of quantum information (Masada et al., 2015).
@@
Acknowledgments
We thank Amir Abolfazli, Timko Dubielzig, Ludwig Krinner, Aditya Mohan, and Kilian Stahl for helpful discussions, Pratik Chakraborty and Manuel Schimanski for help in the lab, and Nived Johny, Jonas Junker, Arindam Saha, and Bernd Schulte for both. LR thanks Mika Nina Seibicke and Alina Jankowski for lab sitting and emotional support.
This work was supported by the Quantum Valley Lower Saxony (QVLS), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through SFB 1227 (DQ-mat), Germany’s Excellence Strategy EXC-2123 QuantumFrontiers 390837967, and Excellence Strategy EXC-2122 PhoenixD 390833453, the BMBF projects ATIQ and QuBRA, the BMWK project ProvideQ, and the Australian Research Council Centre of Excellence for Quantum Computation and Communication Technology (project number CE170100012).
@@ Appendix A Additional details about the experimental setup
Our setup includes the following components: We use a $1064\text{\,}\mathrm{nm}$ laser (Mephisto, Coherent), a single mode polarization-maintaining fiber, and a Schäfter+Kirchhoff fiber collimator (60FC-SF-4-M8-08) at the input side of the fiber. The measurements of laser power are done with power meters (Thorlabs PM160, measurement error $1\%$). In front of the experiment, we place a partially reflecting mirror to measure a fraction of the laser light with an additional power meter for power reference. The measurement is used to pause training in the event of a laser failure and to track power fluctuations to determine the maximum power level. In this way, we can determine the coupling efficiency with an error of $2\%$. The input power is set to $1.00(1)\text{\,}\mathrm{mW}$.
For training on the experiment, we used an NVIDIA GeForce RTX 4070 GPU. In addition to the usual packages (Harris et al., 2020; Abadi et al., 2015; Waskom, 2021; Hunter, 2007), we used PyLabLib, Thorlabs Kinesis, PyVisa, Keysight Connection Expert and safe-exit (Shkarin, 2024; Thorlabs, ; Grecco et al., 2023; Keysight, ; JianAo, ) for communication with the experiment.
@@ Appendix B Explanation of the imprecision in the mirror steering motors and its characterization
Backlash is a phenomenon that is present when a load is not directly connected to a motor, such as in geared mechanical systems (Nordin & Gutman, 2002). Dependent on the exact geometry of the system, i.e., mechanical tolerances, amount of gears, etc., it may resemble hysteresis between the expected and actual position or a dead-zone, where moving the actuator has no effect on the actual position whenever the rotational direction is reversed. It is thus hard to model and predict a priori. Control has to be implemented based on the specific system and its use. These control schemes include hysteresis models, dead-zone models, and PI control. However, additional sensors are needed to get accurate feedback if multiple actuators are used. Step loss results from the difference in static and dynamic torque of a motor. The motor steps result in a linear actuation, which changes the tilt of the mirror mount. Different tilt angles lead to different static loading of the motor and gearbox. This may lead to the initial step(s) being lost, as the motor can not deliver the starting torque, resulting in a partial step. Without feedback from, e.g., an encoder, this leads to a difference between the expected and actual position. Lastly, the non-orthogonal degrees of freedom are a result of the kinematic mirror mounts used and their mounting. Usually, this error is small for well-designed kinematic mirror mounts.
@@ Appendix C Environment and agent tuning on virtual testbed
$P(x_{m1},y_{m1},x_{m2},y_{m2})=0.92\mathrm{exp}\Bigg{(}-\frac{1}{2}\Bigg{(}$ $\left(\frac{x_{m1}-5470785}{11994}\right)^{2}+\left(\frac{y_{m1}-5573194}{19145}\right)^{2}$
@@ C.1 Reward Hyperparameters
$\times\{1,5,10\}$ $\times\{1,5\}^{2}$ $\times\{1,5\}^{2}$
@@ C.2 Episode Length
We want to find a good trade-off between reaching the goal quickly and reaching it reliably. Using the same parameters as for reward shaping, we tested different episode lengths, in particular, T=5,10,,50T=5,10,...,50. The results are shown in Figure 7.
@@ C.5 Observation
We tested history lengths of n=1,,6n=1,...,6 and the maximum sensible length $n=T=30$ with $P_{\text{goal}}=0.85$. The results can be found in Figure 10 (a).
Figure 10: Comparison of different observations and maximum actions. (a)-(c) show the normalized return in dependence of the training step using the parameters in Tables 1 and 2, TQC, and $P_{\text{goal}}=0.85$ for different history lengths $n$ in (a), leaving out different parts of the observation in (b), and with or without including the absolute positions in the observation in (c). The models with and without the absolute positions were then tested $100$ times each in an environment in which k=0,,4k=0,...,4 means of the underlying Gaussian $\mu_{i}$ were shifted by $\pm\sigma_{i}$, i.e., $\mu_{i}^{\prime}=\mu_{i}\pm\sigma_{i}$. (d) shows the probability of reaching the goal against the number of shifts $k$. The plots show the mean with $2\sigma$ error bands/bars created by multiple runs and, in (a)-(c), smoothing.
@@
Additionally, we tested how agents perform in an environment that includes the absolute position of the actuators in the observation compared to one that does not. The normalized return against the training step can be found for both configurations, using the parameters in Tables 1 and 2, TQC, and $P_{\text{goal}}=0.85$, in Figure 10 (c). As expected, the agent learns faster and reaches a higher normalized return if those absolute positions are included. However, the main application for our agent is to recouple the light into the fiber after other parts of the experiment have been misaligned. That means, that the optimal positions, i.e., the means $\mu_{i}$ of the underlying Gaussian, change, and the agent still has to be able to reach the goal. Hence, we tested the agent $100$ times each in environments in which k=0,,4k=0,...,4 means of the underlying Gaussian $\mu_{i}$ were shifted by $\pm\sigma_{i}$, i.e., $\mu_{i}^{\prime}=\mu_{i}\pm\sigma_{i}$. Figure 10 (d) shows the probability of reaching the goal against $k$. If no shifts occur, the agent with the absolute position as part of the observation performs slightly better than the one without. However, this quickly changes as more shifts are applied. The agent with absolute position performs much worse if any shifts appear. On the other hand, the agent not observing the absolute position performs well independent of shifts.